OpenStudy (anonymous):

a length of a rectangle is 5 inches more than the width. the area of the rectangle is 84 sq. inches. find the width of the rectangle a length of a rectangle is 5 inches more than the width. the area of the rectangle is 84 sq. inches. find the width of the rectangle @Mathematics

OpenStudy (saifoo.khan):

x(x+5) = 84 solve.

OpenStudy (anonymous):

so that would be 2x+5x = 84 right? or is it x^2+5x=84?

OpenStudy (saifoo.khan):

hmm.

OpenStudy (anonymous):

?